OpenStudy (anonymous):

What is the area under the curve y=x^(-3/2) in the first quadrant between x=1 and x=+∞?

OpenStudy (anonymous):

\[\lim_{c \rightarrow \infty}\int\limits_{1}^{c}x^{-3/2}dx=\lim_{c \rightarrow \infty}-2x^{-1/2}|^c_1=\lim_{c \rightarrow \infty}\frac{-2}{c^{1/2}}+2=0+2=2\]

OpenStudy (anonymous):

Thenk you!